The perimeter of the rectangle is 46m. If the width were doubled and the length were increased by 5m, the perimeter would be 66m. What is the length of the original rectangle?
Answer by lwsshak3(11628)

The perimeter of the rectangle is 46m. If the width were doubled and the length were increased by 5m, the perimeter would be 66m. What is the length of the original rectangle?
**
let x=length of original triangle
let y=width of original triangle
2y+2x=perimeter=46
x+y=23
y=23-x
..
2y+x+5=66
2(23-x)+x+5=66
46-2x+x+5=66
x=15
y=23-15=8
length of original triangle=15 m
width of original triangle=8 m